### v1

@@ 4.2 JustGRPO
Standard GRPO assumes a policy $\pi(o_{t}|o_{<t},q)$ accepting a partial sequence $o_{<t}$ with all tokens observed and predicts one token $o_{t}$ at a time, where $q$ is the query. Diffusion language models, however, are architected as sequence-level denoisers that accept a full sequence with mixed observed and masked tokens and predict the original values for all masked positions simultaneously.
By forgoing arbitrary-order generation, we are able to bridge the above gap and rigorously define an AR policy $\pi_{\theta}^{\text{AR}}$ for dLLMs. To obtain the probability of the next token $o_{t}$ given history $o_{<t}$, we construct an input state $\tilde{x}_{t}$ where the past is observed and the future is masked:
$\tilde{x}_{t}=[\underbrace{o_{1},\dots,o_{t-1}}_{\text{Observed}},\underbrace{\texttt{[MASK]},\dots,\texttt{[MASK]}}_{\text{Masked}}].$ (3)
Although the diffusion language model outputs predictions for all masked positions, the autoregressive policy concerns only the next token $o_{t}$. We thus define $\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)$ as the probability distribution of $o_{t}$:
$\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q)\triangleq\text{Softmax}(f_{\theta}(\tilde{x}_{t}))_{t},$ (4)
where $f_{\theta}$ denotes the model logits. Consequently, the likelihood of a complete reasoning chain $o$ is exactly computable as:
$\pi_{\theta}^{\text{AR}}(o|q)=\prod_{t=1}^{L}\pi_{\theta}^{\text{AR}}(o_{t}|o_{<t},q).$ (5)
@@
The above formulation enables the direct application of standard GRPO to diffusion language models. For each query $q$, we sample a group of outputs $\{o^{i}\}_{i=1}^{G}$ using the old policy $\pi_{\theta_{\text{old}}}^{\text{AR}}$. The objective is:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\text{AR}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o^{i}|}\sum_{t=1}^{|o^{i}|}\left(\min\left(\rho_{i,t}\hat{A}_{i,t},\text{clip}(\rho_{i,t},1-\varepsilon,1+\varepsilon)\hat{A}_{i,t}\right)-\beta\mathbb{D}_{\text{KL}}\right)\right],$ (6)
where $\rho_{i,t}=\frac{\pi_{\theta}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}{\pi_{\theta_{\text{old}}}^{\text{AR}}(o^{i}_{t}\mid o^{i}_{<t},q)}$ is the probability ratio between the current and old policies and $\hat{A}_{i,t}$ denotes the group-standardized advantage.
@@
One might worry whether training in AR mode turns the diffusion language model into a standard autoregressive model. This is not the case. The AR constraint is applied only during training to correctly assign credit. It refines the model’s joint distribution $p(o)$ without altering the underlying architecture. At inference time, the model retains its conditional independence properties, still allowing us to employ parallel samplers Ben-Hamu et al. (2025) to accelerate decoding. JustGRPO thus achieves the reasoning depth of autoregressive models while preserving the inference speed of dLLMs (see Section 5.2).
@@ 5 Experiments
We apply JustGRPO on LLaDA-Instruct Nie et al. (2025) and evaluate its effectiveness on four standard reasoning and coding benchmarks: GSM8K, MATH, HumanEval, and MBPP. For mathematical tasks, we train on the official training split of each dataset following Zhao et al. (2025); Ou et al. (2025). For coding tasks, we train on a subset of AceCoder-87K Zeng et al. (2025) following Gong et al. (2025); Ou et al. (2025). We apply JustGRPO directly to the models without additional task-specific SFT. Our training recipe largely follows Huang et al. (2025b). To evaluate the trained models, we follow the standard LLaDA evaluation protocol Nie et al. (2025), which applies low-confidence remasking together with semi-autoregressive decoding in blocks of 32 tokens, using a sampling temperature of 0. We evaluate all benchmarks at generation lengths of 128, 256, and 512 following Zhao et al. (2025); Ou et al. (2025). More details on training are provided in Appendix A.
@@ 5.1 Main Results
Table 1 reports the system-level comparison. We observe that simplifying the training objective to a standard autoregressive formulation yields consistent improvements over methods specifically designed for dLLMs.
@@
On GSM8K, JustGRPO achieves 89.1% accuracy (seq len 256), improving upon the previous best method, SPG, by a non-trivial margin (3.0%). This trend generalizes to the more challenging MATH-500 benchmark, where our approach outperforms ESPO by 6.1%. These results challenge the prevailing assumption that RL for diffusion models requires optimizing over the full combinatorial space of denoising trajectories. Instead, treating the dLLM as a sequential generator during training appears more effective for credit assignment in logic-heavy tasks.
@@
A potential concern is that AR constraints might overfit the model to specific trajectory lengths. However, we observe robust performance gains across varying sequence lengths (128, 256, 512). This stability suggests that the policy has improved its fundamental reasoning capability, i.e., the ability to navigate logical branches, rather than merely memorizing fixed-length patterns.
Table 1: System-level comparison. RL post-training approaches on LLaDA-Instruct. JustGRPO consistently achieves state-of-the-art performance across all tasks and sequence lengths. LLaDA-1.5 and LLADOU are excluded from the comparison as LLaDA-1.5 is trained on a privately collected dataset at a significantly larger scale, while LLADOU modifies the base architecture with an auxiliary module.
@@
LLaDA-1.5 - 83.3 - - - - 29.3 39.6 51.9 39.6 39.9 38.8
LLaDOU - 88.1 - - 41.1 - - 59.1 - - 51.6 -
D1 73.2 81.1 82.1 33.8 38.6 40.2 - - - - - -
WD1 77.2 80.8 82.3 33.3 34.4 39.0 - - - - - -
$d$-TreeRPO - 81.2 82.6 - 37.7 38.9 - - - - - -
ESPO 80.0 82.3 83.7 36.0 39.0 43.4 28.1 42.1 50.0 47.4 44.6 44.2
GDPO 78.4 82.8 84.5 33.2 39.6 41.4 26.2 39.6 39.0 43.6 50.6 47.1
SPG 78.5 86.1 84.5 33.4 40.0 41.8 - - - - - -
JustGRPO (Ours) 83.8 89.1 89.8 39.0 45.1 45.2 37.8 49.4 48.7 50.6 52.4 49.0
@@ 5.2 JustGRPO Preserves Parallel Decoding
Figure 7: JustGRPO preserves the parallel decoding capability of dLLMs. Interestingly, when compared to original instruct model, accuracy gains are larger with more parallel tokens, likely due to more robust reasoning capabilities after JustGRPO training. We adopt training-free EB-sampler Ben-Hamu et al. (2025) for parallel decoding.
We further investigate whether the AR training constraint compromises the model’s inherent diffusion capabilities. We employ the training-free Entropy Bounded (EB) Sampler Ben-Hamu et al. (2025) to evaluate inference performance under varying degrees of parallelism (tokens per step).
Figure 7 reveals that our model not only retains full compatibility with parallel decoding but exhibits a strictly superior trade-off between speed and accuracy. Surprisingly, the performance gain becomes more pronounced as parallelism increases. As shown in the MBPP and HumanEval results, while the baseline’s performance degrades sharply with more aggressive parallel steps, our model maintains stability. For instance, on MBPP, the accuracy gap expands from +10.6% at conservative settings (1 token/step) to +25.5% at aggressive settings ($\sim$5 tokens/step).
This observation suggests that JustGRPO does not merely fit a specific decoding path; rather, it learns a more robust reasoning manifold that is resilient to the approximation errors inherent in parallel sampling. The AR formulation thus acts as an effective training scaffold that refines the joint distribution $p(o)$ during optimization, providing a more stable foundation for parallel samplers to exploit at inference.
@@ 6 Related Work
While early studies validated the utility of arbitrary order generation in constrained tasks like Sudoku and Zebra Puzzles Ye et al. (2024); Nie et al. (2025); Kim et al. (2025), recent research has begun to investigate its value in standard reasoning domains. One line of research, exemplified by Dream-Coder Xie et al. (2025) and DiffuCoder Gong et al. (2025), shows that diffusion models can naturally exhibit non-standard, human-like decoding behaviors (e.g., sketch-first reasoning) without any explicit supervision on generation order. DiffuCoder Gong et al. (2025) further quantifies this phenomenon by demonstrating that higher sampling temperatures reduce AR-ness, with increased order randomness correlating with improved output diversity. Another line of research, exemplified by P2 Peng et al. (2025) and LLaDOU Huang et al. (2025b), explicitly optimizes the decoding strategy for better generation order, thus better performance. Despite these advances, it remains unclear whether the observed gains primarily arise from better exploitation of existing solution patterns encoded in the data and model, or whether order arbitrariness itself enables qualitatively new reasoning strategies that are unattainable under a purely autoregressive decoding regime.
@@ 7 Conclusion
Therefore, eliciting the reasoning capability of dLLMs can be simpler. By operating dLLMs in a standard autoregressive manner, we enable the direct application of Group Relative Policy Optimization (GRPO) without any complex adaptations tailored for order arbitrariness. This intentional constraint paradoxically yields a significant upgrade in reasoning performance, while fully preserving the parallel decoding capabilities of dLLMs. By returning to the basic, natural left-to-right order of language modeling, we hope to encourage a re-examination of its real value in the training of next-generation diffusion models.
@@ A.3 Reward Function
Appendix B More Analysis Results
To further validate the robustness of our findings in Section 3, we conduct extended analyses on the HumanEval benchmark using the LLaDA-Instruct model Nie et al. (2025). We investigate the impact of temperature, sampling strategies, and decoding block sizes on reasoning performance.
@@ Appendix B More Analysis Results
Figure 9: (a) Sampling algorithm comparison. (b) Correlation between different sampling algorithms and AR in per-problem accuracy. We compare different sampling algorithms in pass@k performance and correlation with AR in per-problem accuracy.
The results show that although more sophisticated sampling algorithms can achieve better pass@k performance than default confidence-based sampling, they still cannot catch up with AR order. Meanwhile, these better pass@k algorithms also show slightly worse pass@1 performance compared to confidence-based sampling, making the overall pass@$k$ curve closer to AR order’s pass@$k$ curve. To investigate this similarity, we calculate the per-problem accuracy correlation between different sampling algorithms and the AR mode (Figure 9(b)). We observe that algorithms with higher scaling potential (higher pass@128) consistently show stronger correlation with AR, with the most effective method (Neg-Entropy) achieving a correlation of $0.970$. This suggests that sampling algorithms with better pass@$k$ tend to behave more like AR in terms of task-level performance characteristics.
@@
B.4 Entropy Comparison Results on More Forking Tokens
@@
B.5 Training Efficiency Analysis
@@
Figure 12 compares JustGRPO against ESPO Ou et al. (2025), a representative approximation-based baseline. It can be observed that ESPO saturates early, effectively plateauing around 10 hours. In contrast, JustGRPO exhibits a superior efficiency frontier: it surpasses the baseline’s peak accuracy in less time and maintains a continuous scaling trend. Furthermore, motivated by the observation that reasoning is primarily steered by a subset of high-entropy forking tokens Wang et al. (2025c), we demonstrate that JustGRPO allows for further acceleration via a simple heuristic: restricting gradient computation to the top-25% high-entropy tokens (Ours-Fast). This modification yields even faster convergence without compromising the final performance. We prioritize simplicity in the current implementation and leave further efficiency optimizations for future work.

### v2

@@ 4.2 JustGRPO
Standard GRPO assumes a policy $\pi(o_{k}|o_{<k},q)$ accepting a partial sequence $o_{<k}$ with all tokens observed and predicts one token $o_{k}$ at a time, where $q$ is the query. Diffusion language models, however, are architected as sequence-level denoisers that accept a full sequence with mixed observed and masked tokens and predict the original values for all masked positions simultaneously.
By forgoing arbitrary-order generation, we are able to bridge the above gap and rigorously define an AR policy $\pi_{\theta}^{\text{AR}}$ for dLLMs. To obtain the probability of the next token $o_{k}$ given history $o_{<k}$, we construct $\tilde{x}_{k}$ where the past is observed and the future is masked:
$\tilde{x}_{k}=[\underbrace{o_{1},\dots,o_{k-1}}_{\text{Observed}},\underbrace{\texttt{[MASK]},\dots,\texttt{[MASK]}}_{\text{Masked}}].$ (5)
Although the diffusion language model outputs predictions for all masked positions, the autoregressive policy concerns only the next token $o_{k}$. We define $\pi_{\theta}^{\text{AR}}(\cdot|o_{<k},q)$ as follows:
$\pi_{\theta}^{\text{AR}}(\cdot|o_{<k},q)\triangleq\text{Softmax}(f_{\theta,k}(\tilde{x}_{k},q)),$ (6)
where $f_{\theta,k}$ denotes the model logits at position $k$. Consequently, the likelihood of a complete reasoning chain $o$ is exactly computable as:
$\pi_{\theta}^{\text{AR}}(o|q)=\prod_{k=1}^{|o|}\pi_{\theta}^{\text{AR}}(o_{k}|o_{<k},q).$ (7)
@@
The above formulation enables the direct application of standard GRPO to diffusion language models. For each query $q$, we sample a group of outputs $\{o_{i}\}_{i=1}^{G}$ using the old policy $\pi_{\theta_{\text{old}}}^{\text{AR}}$. The objective is:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\text{AR}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\left(\min\left(\rho_{i,k}\hat{A}_{i,k},\text{clip}(\rho_{i,k},1-\varepsilon,1+\varepsilon)\hat{A}_{i,k}\right)-\beta\mathbb{D}_{\text{KL}}\right)\right],$ (8)
where $\rho_{i,k}=\frac{\pi_{\theta}^{\text{AR}}(o_{i,k}\mid o_{i,<k},q)}{\pi_{\theta_{\text{old}}}^{\text{AR}}(o_{i,k}\mid o_{i,<k},q)}$ is the probability ratio between the current and old policies and $\hat{A}_{i,k}$ denotes the group-standardized advantage.
@@
One might worry whether training in AR mode turns the diffusion language model into a standard autoregressive model. This is not the case. The AR constraint is applied only during training to correctly assign credit. It refines the model’s joint distribution without altering the underlying architecture. At inference time, the model retains its conditional independence properties, still allowing us to employ parallel samplers Ben-Hamu et al. (2025) to accelerate decoding. JustGRPO thus benefits from reasoning exploration of AR order while preserving the inference speed of dLLMs (see Section 5.2).
@@ 5 Experiments
We apply JustGRPO on LLaDA-Instruct Nie et al. (2025) and evaluate its effectiveness on four standard reasoning and coding benchmarks: GSM8K, MATH-500, HumanEval, and MBPP. For mathematical tasks, we train on the official training split of each dataset following Zhao et al. (2025); Ou et al. (2025). For coding tasks, we train on a subset of AceCoder-87K Zeng et al. (2025) following Gong et al. (2025); Ou et al. (2025). We apply JustGRPO directly to the models without additional task-specific SFT. Our training recipe largely follows Huang et al. (2025b). To evaluate the trained models, we follow the standard LLaDA evaluation protocol Nie et al. (2025), which applies low-confidence remasking together with semi-autoregressive decoding in blocks of 32 tokens, using a sampling temperature of 0. We evaluate all benchmarks at generation lengths of 128, 256, and 512 following Zhao et al. (2025); Ou et al. (2025). More details on training are provided in Appendix A.
@@ 5.1 Main Results
Table 1 reports the system-level comparison. We observe that simplifying the RL objective to a standard GRPO formulation yields consistent improvements over methods specifically designed for dLLMs.
@@
On GSM8K, JustGRPO achieves 89.1% accuracy (seq len 256), improving upon the previous best method, SPG, by a non-trivial margin (3.0%). This trend generalizes to the more challenging MATH-500 benchmark, where our approach outperforms ESPO by 6.1%. These results challenge the prevailing assumption that RL for diffusion models requires optimizing over the full combinatorial space of denoising trajectories. Instead, treating the dLLM as a sequential generator during training appears more effective for credit assignment in general reasoning tasks.
@@
A potential concern is that AR constraints might overfit the model to specific trajectory lengths. However, we observe robust performance gains across varying sequence lengths (128, 256, 512). This stability suggests that the policy has improved its underlying reasoning capability rather than merely memorizing fixed-length patterns.
Table 1: System-level comparison of RL post-training approaches on LLaDA-Instruct. JustGRPO achieves better performance than existing diffusion-specific RL methods. LLaDA-1.5 and LLADOU are excluded from the comparison as LLaDA-1.5 is trained on a privately collected dataset at a significantly larger scale, while LLaDOU modifies the base architecture with an auxiliary module.
@@
LLaDA-1.5 Zhu et al. (2025) - 83.3 - - - - 29.3 39.6 51.9 39.6 39.9 38.8
LLaDOU Huang et al. (2025b) - 88.1 - - 41.1 - - 59.1 - - 51.6 -
d1 Zhao et al. (2025) 73.2 81.1 82.1 33.8 38.6 40.2 - - - - - -
wd1 Tang et al. (2025) 77.2 80.8 82.3 33.3 34.4 39.0 - - - - - -
$d$-TreeRPO Pan et al. (2025) - 81.2 82.6 - 37.7 38.9 - - - - - -
ESPO Ou et al. (2025) 80.0 82.3 83.7 36.0 39.0 43.4 28.1 42.1 50.0 47.4 44.6 44.2
GDPO Rojas et al. (2025) 78.4 82.8 84.5 33.2 39.6 41.4 26.2 39.6 39.0 43.6 50.6 47.1
SPG Wang et al. (2025a) 78.5 86.1 84.5 33.4 40.0 41.8 - - - - - -
JustGRPO 83.8 89.1 89.8 39.0 45.1 45.2 37.8 49.4 48.7 50.6 52.4 49.0
@@ 5.2 JustGRPO Preserves Parallel Decoding
Figure 7: JustGRPO preserves the parallel decoding capability of dLLMs. Interestingly, when compared to original instruct model, accuracy gains are larger with more parallel tokens, likely due to more robust reasoning capabilities after JustGRPO training. We adopt training-free EB-sampler Ben-Hamu et al. (2025) for parallel decoding. Generation length is set to 256.
We further investigate whether the AR training compromises dLLMs’ parallel decoding capabilities. We employ the training-free Entropy Bounded (EB) Sampler Ben-Hamu et al. (2025) to evaluate inference performance under varying degrees of parallelism (tokens per step).
Figure 7 shows that our model retains full compatibility with parallel decoding and exhibits a strictly superior trade-off between speed and accuracy than the original LLaDA-Instruct model. Interestingly, the performance gain becomes more pronounced as parallelism increases. While the baseline’s performance degrades sharply with high parallel steps, our model maintains relatively more stable. For instance, on MBPP, the accuracy gap expands from +10.6% at conservative settings (1 token/step) to +25.5% at aggressive settings ($\sim$5 tokens/step).
This observation suggests that JustGRPO does not merely fit a specific decoding path; rather, it learns a more robust reasoning manifold that is resilient to the approximation errors inherent in parallel sampling. In other words, the AR formulation acts as a training “scaffold” that refines the joint distribution $p(o)$ more effectively, providing a more stable foundation for inference.
@@ 6 Related Work
While early studies validated the utility of arbitrary order generation in constrained tasks like Sudoku and Zebra Puzzles Ye et al. (2024); Nie et al. (2025); Kim et al. (2025), recent research has begun to investigate its value in general reasoning domains. One line of research, exemplified by Dream-Coder Xie et al. (2025) and DiffuCoder Gong et al. (2025), shows that diffusion models can naturally exhibit non-standard, human-like decoding behaviors (e.g., sketch-first reasoning) without any explicit supervision on generation order. DiffuCoder Gong et al. (2025) further quantifies this phenomenon by demonstrating that higher sampling temperatures reduce AR-ness, with increased order randomness correlating with improved output diversity. Another line of research, exemplified by P2 Peng et al. (2025) and LLaDOU Huang et al. (2025b), explicitly optimizes the decoding strategy for better generation order, thus better performance. Despite these advances, it remains unclear whether the observed gains primarily arise from better exploitation of existing solution patterns encoded in the data and model, or whether order arbitrariness itself enables qualitatively new reasoning strategies that are unattainable under a purely autoregressive decoding regime.
@@ 7 Conclusion
Therefore, the elicitation of reasoning capability for dLLMs can be simpler. By operating dLLMs in a standard autoregressive manner, we enable the direct application of Group Relative Policy Optimization (GRPO) without any complex adaptations tailored for order arbitrariness. This intentional constraint paradoxically yields a significant improvement in reasoning performance, while fully preserving the parallel decoding capabilities of dLLMs. By returning to the basic, natural left-to-right order of language modeling, we hope to encourage a re-examination of its real value in the development of next-generation language models.
@@ A.3 Reward Function
Appendix B More Analysis Results on Reasoning Boundary
To further validate the robustness of our findings about reasoning boundary in Section 3, we conduct extended analyses on the HumanEval benchmark using the LLaDA-Instruct model Nie et al. (2025). We investigate the impact of temperature, sampling strategies, and decoding block sizes on reasoning performance.
@@ Appendix B More Analysis Results on Reasoning Boundary
Figure 9: (a) Sampling algorithm comparison. (b) Correlation between different sampling algorithms and AR in per-problem accuracy. Algorithms with higher Pass@$k$ performance tend to show higher correlation with AR in problem-level performance.
The results show that although more sophisticated sampling algorithms can achieve better Pass@$k$ performance than default confidence-based sampling, they still cannot catch up with AR order. Meanwhile, these better Pass@$k$ algorithms also show slightly worse Pass@$1$ performance compared to confidence-based sampling, making the overall Pass@$k$ curve closer to AR order’s Pass@$k$ curve. To investigate this similarity, we calculate the per-problem accuracy correlation between different sampling algorithms and the AR mode (Figure 9(b)). We observe that algorithms with higher scaling potential (higher Pass@$128$) consistently show stronger correlation with AR, with the most effective method (Neg-Entropy) achieving a correlation of $0.970$. This suggests that sampling algorithms with better Pass@$k$ tend to behave more like AR in performance characteristics.
@@
Appendix C More Results on Entropy Degradation
@@
Appendix D Training Efficiency Analysis of JustGRPO
@@
Figure 12 compares JustGRPO against ESPO Ou et al. (2025), a representative approximation-based baseline. It can be observed that ESPO saturates earlier, effectively plateauing around 10 hours. In contrast, JustGRPO exhibits a superior efficiency frontier: it surpasses the baseline’s peak accuracy in less time and maintains a continuous scaling trend. Furthermore, motivated by the observation that reasoning is primarily steered by a subset of high-entropy forking tokens Wang et al. (2025c), we demonstrate that JustGRPO allows for further acceleration via a simple heuristic: restricting gradient computation to the top-25% high-entropy tokens (Ours-Fast). This modification yields faster convergence and even slightly better performance. We prioritize simplicity in the current implementation and leave further efficiency optimizations for future work.